A particle of mass π‘šπ‘š collides elastically with a particle of mass 3π‘šπ‘š which is at rest. In the




𝐢𝐢 βˆ’frame, the particle of mass π‘šπ‘š moves with speed 3 π‘šπ‘š/𝑠𝑠 at an angle of 60π‘œπ‘œ after the




collision. Find the final velocity for the mass π‘šπ‘š in the laboratory frame.

1
Expert's answer
2022-01-05T14:30:16-0500

Given quantities:

"m_1=m"

"m_2=3m"

"v_1=" "3 \frac{m}{s}"

"\alpha=60^o"

"v_2=0"

From conservation of momentum

"P_1+P_2=P_1^{'}+P_2^{'}"

"m_1v_1+m_2v_2=(m_1+m_2)U"

"mv_1cos60^o+o=4mU"

"U=\frac{3*0.5}{4}=\frac{3}{8}=0.375\frac{m}{s}"